OpenStudy (anonymous):

Chris, who is 6 feet tall, is walking away from a street light pole 30 feet high at a rate of 2 feet per second. 1. How fast is his shadow increasing in length when he is 24 feet from the pole? 2. How fast is the tip of his shadow moving when he is 24 feet from the pole? @Physics

OpenStudy (amistre64):

im learning to read these better :)

OpenStudy (amistre64):



OpenStudy (amistre64):

we can relate the whole thing with similar triangles

OpenStudy (amistre64):

\[\frac{30}{w+s}=\frac{6}{s}\] \[30s=6(w+s)\] derive it with respect to time \[s'=6(w'+s')\] \[s'=6w'+6s'\] \[s'-6s'=6w'\] \[-5s'=6w'\] \[s'=\frac{6w'}{-5}\] did i get that right?

OpenStudy (amistre64):

since hes walking, w' = 2 but i gotta dbl chk the calculations i made

OpenStudy (amistre64):

\[30s=6(w+s)\] \[30s=6w+6s\] \[30s-6s=6w\] \[24s=6w\] \[s=\frac{1}{4}w\] \[s'=\frac{1}{4}w'\]

OpenStudy (amistre64):

im missing something i can feel it

OpenStudy (amistre64):

i dropped the 30 by mistake the first time, but we should have a w or s in there someplace otherwise the rate is the same regardless of distances

OpenStudy (anonymous):

so s'=1/4w', which is 1/4(2)=.5 ft/s?

OpenStudy (amistre64):

yes, so far as that goes, but still ... im concerned

OpenStudy (amistre64):

the shadow growth is s'; the speed at which it moves is w'+s' tho

OpenStudy (anonymous):

so 2.5 ft/s?

OpenStudy (amistre64):

yeah, as long as the calculations are good

OpenStudy (anonymous):

I see, thank you that was a huge help. What about 2. How fast is the tip of his shadow moving when he is 24 feet from the pole?

OpenStudy (amistre64):

let me try this from another vantage point: \[\frac{30}{6}=\frac{(w+s)}{s}\]derived we get: \[0=s(w'+s')-s'(w+s)\] \[0=sw'+ss'-s'w+ss'\] \[0=sw'-s'w\] \[s'=\frac{sw'}{w}\] \[s'=\frac{2s}{24}\] \[\frac{30}{6}=\frac{(24+s)}{s}\] \[30s=6(24+s)\] \[30s=144+6s;s=144/24=6\] \[s'=\frac{2(6)}{24}=\frac{1}{2}ft/sec\]is how fast the shadow is growing add 2 feet per sec to that from teh walking speed to calculate the tips speed: 5/2 ft/sec

OpenStudy (anonymous):

thank you soo much!

OpenStudy (anonymous):

Sorry, one more question: How fast is the tip of his shadow moving when he is 24 feet from the pole? I don't know what I should do?